OpenStudy (1018):

what is the equation of the sine cruve with the period 20pi and amplitude 10?

OpenStudy (1018):

my answer : \[10\sin(\frac{ x }{ 10 })\]

OpenStudy (1018):

is this correct?

OpenStudy (michele_laino):

yes, it is correct. In general we have this formula: \[y\left( x \right) = A\sin \left( {\frac{{2\pi x}}{l}} \right)\] being \(l\) the new period

OpenStudy (1018):

Thank you!

OpenStudy (michele_laino):

:)